OpenStudy (anonymous):

7 buds make a cigarette you have 49 buds how many cigarettes can you make?

OpenStudy (anonymous):

7 buds can make .................1 cigarette 1 buds can make.....................1/7 cigarette 49 buds can make .....................1/7(49)cigrattes = 1/7 x 49= 7 ans

OpenStudy (anonymous):

no shrishti..try again..

OpenStudy (anonymous):

Well it looks like 7 doesn't it but it is actually 8. Think about theb uds made frome the 7 cigarette constructed from the original 49. They will all produce a bud allowing one more cigarette to be made.

OpenStudy (anonymous):

yes it is 8

OpenStudy (anonymous):

amazing answer..

OpenStudy (anonymous):

good gianfranco

OpenStudy (nowhereman):

Smoking will kill you, so I would not make any!

OpenStudy (anonymous):

good gianfranco and dipankar

OpenStudy (anonymous):

No , its not 8 , its 7 only. we have asked how many cigratte from 49 buds we have given 7 bud make 1 cigratte and 49 can make 7 only not 8

OpenStudy (anonymous):

Shrishti, the 7 you make will create 7 buds which will in turn make another cigarette.

OpenStudy (anonymous):

shrishti..when you smoke a cigarette..a bud is left..so think laterally

OpenStudy (anonymous):

hmmnnnnnn